# Compression Hacking: A Supplementary Perspective on Informatics Properties of Language Models from Geometric Distortion

### v2

@@ Abstract
Recently, the concept of “compression as intelligence” has provided a novel informatics metric perspective for language models (LMs), emphasizing that highly structured representations signify the intelligence level of LMs. However, from a geometric standpoint, the representation space of highly compressed LMs tends to degenerate into a highly anisotropic state, which hinders the LM’s ability to comprehend instructions and directly impacts its performance. We found this compression-anisotropy synchronicity is essentially the “Compression Hacking” in LM representations, where noise-dominated directions tend to create the illusion of high compression rates by sacrificing spatial uniformity. Based on this, we propose three refined compression metrics by incorporating geometric distortion analysis and integrate them into a self-evaluation pipeline. The refined metrics exhibit strong alignment with the LM’s comprehensive capabilities, achieving Spearman correlation coefficients above 0.9, significantly outperforming both the original compression and other internal structure-based metrics. This confirms that compression hacking substantially enhances the informatics interpretation of LMs by incorporating geometric distortion of representations.
Compression Hacking: A Supplementary Perspective on Informatics Properties of Language Models from Geometric Distortion
Jianxiang Zang1, Meiling Ning2, Yongda Wei3, Shihan Dou1, Jiazheng Zhang1, Nijia Mo4, Binhong Li5, Tao Gui1thanks: Corresponding author., Qi Zhang1, Xuanjing Huang1∗ 1 Fudan University, 2 Beijing University of Posts and Telecommunications, 3 George Mason University 4 Shanghai University of International Business and Economics 5 Hong Kong University of Science and Technology (Guangzhou) jxzang25@m.fudan.edu.cn, tgui@fudan.edu.cn
@@ 1 Introduction
Recently, significant efforts have been devoted to exploring the mechanisms by which language models (LMs) process information internally, driving the development of LM self-evaluation Wei et al. (2024); Wang et al. (2024a, b) independent of specific tasks and model outputs. The concept of “compression as intelligence” Sutskever (2023); Deletang et al. (2023); Chen et al. (2025) has provided a novel informatics interpretation for LMs, emphasizing that LMs eliminate redundant information through training while their representation spaces typically evolve from disordered to structured states. This property leads to a compression-based evaluation metric for LMs that utilizes differential entropy of representations, aiming to reflect model capabilities with their internal structural organization Pichler et al. (2022); Zhouyin and Liu (2023); Li et al. (2025). Existing studies have demonstrated strong alignment between this metric and LM scale Wei et al. (2024); Li et al. (2025), which we have also empirically validated. However, as evidenced by the intuitive case where 175B GPT-3 Brown et al. (2020) exhibits inferior overall capabilities compared to 32B Qwen2.5-Instruct Hui et al. (2024), compression from a purely informatics standpoint, cannot fully align with LM capabilities, especially when comparing models from different families. Therefore, our research motivation is: Beyond information compression, what other properties should a metric quantify to effectively interpret the LMs’ intelligence level, and how should we model the relationships between these properties?
Relevant studies have shown that differences in model architecture and training paradigms inevitably lead to variations in the geometric structure of representations Mimno and Thompson (2017); Gao et al. (2019a); Skean et al. (2025). From a geometric standpoint, we were surprised to observe that LMs with high information compression tend to exhibit representation spaces that degenerate into highly anisotropic, distorted states. Highly anisotropic representations indicate varying sensitivity to semantic changes across different dimensions, which can hinder language models’ ability to comprehend instructions and consequently degrade their performance Demeter et al. (2020); Yu et al. (2022); Rudman and Eickhoff (2024).
@@
Using this framework, we evaluate 18 open-source LMs and conduct meta-evaluations on factuality, reasoning, math, and knowledge tasks to obtain ground-truth capability scores. Extensive experiments demonstrate that the refined metrics exhibit strong alignment with the LM’s comprehensive capabilities, achieving Spearman correlation coefficients above 0.9, which significantly outperforms both the original compression and other internal structure-based metrics. This validating the compression hacking substantially enhances the informatics interpretation of LMs by incorporating geometric distortion analysis of representations. The main contributions are summarized as follows:
@@
  • The refined metrics exhibit significantly stronger alignment with LM’s comprehensive capabilities compared to original compression metric, thereby establishing a task-agnostic self-evaluation perspective for LMs. Codes available here.
@@ 2 Compression Hacking
In this section, we analyze the compression-anisotropy synchronicity in LM representations, where highly compressed LMs tend to exhibitin-context representations with strong anisotropy. Our investigation proceeds in two stages: First, we quantify both compression and anisotropy metrics by examining the internal structure of LM representations (covariance matrices). We then fit regression curves to model the relationship between anisotropy and compression, verifying it’s statistical significance. Second, through mechanistic analysis, we identify the underlying cause of this phenomenon, what we term “compression hacking”.
@@ 2.1 Preliminary: Differential Entropy based Compression Metric
The compression perspective provides an information-theoretic foundation for LM evaluation, revealing the intrinsic connections between model scale, generalization capability, and data volume, thus offering theoretical guidance for optimizing model design Pichler et al. (2022); Sutskever (2023); Deletang et al. (2023); Wei et al. (2024); Chen et al. (2025). Related studies have shown that the differential entropy $\mathcal{H}_{\text{DE}}(\mathbf{Z})=-\mathbb{E}_{\bm{w}\sim\mathcal{V}}\mathbf{z}(\bm{w})\log\mathbf{z}(\bm{w})$ of LM representations $\mathbf{z}(\bm{w})$ can reflect their compression capacity Chen et al. (2023a); Zhouyin and Liu (2023); Li et al. (2025). Lower differential entropy suggests that the representations formed by nonlinear transformation, which removes redundant information, are closer to optimal coding. These representations exhibit more concentrated distributions and lower uncertainty, reflecting more efficient information compression Delétang et al. (2023). Semantic Volume leverages this property to model representation uncertainty Li et al. (2025).
We thus define compression metric as the negative differential entropy of representations (i.e., $\mathcal{C}_{\text{DE}}(\mathbf{Z})\overset{\rm def}{=}-\mathcal{H}_{\text{DE}}(\mathbf{Z})$). Since the differential entropy is equivalent to the logdet estimator Chen et al. (2023a) of their covariance matrix, the compression metric follows the definition in Eq. 2.
$\mathcal{C}_{\text{DE}}(\mathbf{Z})$ =$\overset{\rm def}{=}-\frac{1}{2}\text{logdet}\left(\Sigma_{\mathbf{Z}}\right)=-\frac{1}{2}\sum_{d=1}^{D}\log\lambda_{d}$ (2)
Figure 1: Comparison of compression metrics across different models and their corresponding ground-truth comprehensive capabilities, categorized into intra-family and cross-family comparsions.
We first conducted preliminary exploration to assess whether differential entropy-based compression metrics effectively reflect LM capabilities. Our evaluation included both intra-family (OPT family) and cross-family tests (Qwen2.5-3b-Instruct, Qwen2.5-7b, LLaMA3.1-8b, and OPT-13b), with ground-truth settings following Section 4.1. As shown in Figure 1, we found that compression metrics showed only positive correlations with model scale, consistent with related studies Wei et al. (2024); Li et al. (2025). However, Figure 1(left) indicates that differential entropy-based compression is effective only for intra-family evaluation, while Figure 1(right) reveals its limited applicability across diverse architectures and training paradigms. These findings prompted our integration of geometric properties into compression analysis.
@@ 2.2 Anisotropy: The Geometric Property Correlated with Compression
The anisotropy of language models is a geometric property of representations that reflects the non-uniform distribution of semantics across different directions in the representation space Ethayarajh (2019); Cai et al. (2019); Demeter et al. (2020). Highly anisotropic representations hinder LMs’ ability to comprehend instructions Yu et al. (2022); Rudman and Eickhoff (2024), directly impairing their overall capabilities. We performed principal component analysis to visualize thein-context representation spaces of the aforementioned four models. As shown in Figure 2, we made the intriguing observation that models with higher compression levels consistently exhibited greater unevenness in their dimensional distributions, namely, higher anisotropy. This suggests a potential synergistic relationship between compression and anisotropy. If we can quantify this relationship and confirm its statistical significance, it could provide valuable guidance for refining compression metrics.
Current tools for qualitatively and quantitatively analyzing the anisotropy of language models mainly rely on similarity computations of representations Ethayarajh (2019); Cai et al. (2019); Rudman et al. (2022). However, what we need is an anisotropy metric that can establish a connection with entropy-based information compression. Relevant studies Arora et al. (2016); Mu and Viswanath (2018) have shown that the anisotropy measure $\mathcal{A}$ is mathematically defined as formulated in Eq. 3. We aim to extend this measure to relate to the internal structure of representations (eigenvalues of the covariance matrix).
@@
where $\mathcal{Z}(\mathbf{c})=\sum_{\bm{w}\in\mathcal{V}}\exp\left(\mathbf{c}^{\top}\mathbf{z}(\bm{w})\right)$ is the original partition function should approximately be a constant for any unit vector $\mathbf{c}$. $\mathcal{A}$ is a number greater than 1, where larger values indicate stronger anisotropy in the represrntation space. Ideally, this value should be as close to 1 as possible. Considering that $\arg\max_{\|\mathbf{c}\|=1}\mathcal{Z}(\mathbf{c})$ and $\arg\min_{\|\mathbf{c}\|=1}\mathcal{Z}(\mathbf{c})$ do not have closed-form solutions, we attempt to approximate $\mathcal{Z}(\mathbf{c})$ via Taylor expansion as formulated in Eq. 4.
@@ 2.3 Systematic Analysis
Significance Analysis Next, we analyze the significance of compression hacking, which manifests as compression-anisotropy synchronicity. Based on the aforementioned metrics, we calculated the estimates of both compression and anisotropy for instruction representations across four LMs in our preliminary experiments, both of which can be exclusively represented by the eigenvalues of the representation covariance matrix. Given their characteristic patterns, we modeled a linear regression of compression against the logarithmic values of anisotropy, as shown in Figure 3. The regression analysis reveals two key findings through R² and p-values: (1) compression as the dependent variable can be well and significantly explained by anisotropy, and (2) Mann-Whitney U tests McKnight and Najab (2010) confirm statistically significant differences in regression curves across different models.
Figure 3: Regression fitting curves of compression versus anisotropy for different models, along with Mann-Whitney U tests between them. Here, **** denotes statistical significance at the 0.01% levels respectively.
@@ 3.1 Refined Metrics
Spectral Entropy Quantification Figure 2 illustrates that, from the perspective of eigenvalue distribution, the mechanism of compression hacking is that the secondary components causing anisotropy ($\lambda_{d}$) are homologous to the principal components of the compression part ($-\log\lambda_{d}$). Interesting, spectral entropy Roy and Vetterli (2007) precisely models this characteristic, and it is formally equivalent to a compression metric weighted by eigenvalues (Compression (SE)), as formulated in Eq. 8.
@@
Semantic Coefficient of Variation Just as compression-anisotropy synchronicity serves as a distinct manifestation of compression hacking, where compression is characterized by the mean of eigenvalue logarithms (reflecting the overall volume of the embedding space Li et al. (2025)), while anisotropy corresponds to the ratio of extreme eigenvalues (quantifying the variation of semantic embeddings across different dimensions). Thus, we formulate their ratio as the Semantic Coefficient of Variation (Semantic CV) in Eq. 9. This metric accurately characterizes the magnitude of anisotropy relative to information compression in the representation space $\mathbf{Z}$.
@@
Manifold Correction Protocol Numerous studies have proposed train-free “anisotropy razors” to reduce the anisotropy of representation space in a train-free manner, thereby enhancing representational capacity Mu and Viswanath (2018); Su et al. (2021). This inspires us to decouple anisotropy from compression by selecting an appropriate anisotropy razor. Considering the exponential sharp decline in the eigenvalues of principal components corresponding to preceding dimensions due to compression hacking, we propose Principal Component Smoothing (PCS) as an anisotropy razor, inspired by the LW-shrinkage Ledoit and Wolf (2004). By setting a smoothing coefficient $\beta\in[0,1]$ (default value is set to 0.9), we shift the representation space toward principal directions, resulting in a flatter transformed feature spectrum. This transformation is based on the covariance matrix of the representation and is achieved by defining the mapping $\mathcal{T}_{\text{PCS}}$ as formulated in Eq. 10, thereby refining the compression metric Compression (PCS). In Theorem B.2, we prove that under sparse spectrum conditions, the PCS estimator exhibits higher statistical stability than the LW shrinkage.
@@ 4.1 Setup
Models Since our evaluation focuses on the internal structure of model representations, we evaluated 18 open-source language models from three different model families with varying sizes. These families are the LLaMA3 family Grattafiori et al. (2024) (LLaMA3.2-1B, LLaMA3.2-1B-Instruct, LLaMA3.2-3B, LLaMA3.2-3B-Instruct, LLaMA3.1-8B, LLaMA3.1-8B-Instruct), Qwen2.5-Instruct family Hui et al. (2024) (0.5B, 1.5B, 3B, 7B, 14B, 32B), OPT family Zhang et al. (2022a) (0.125B, 1.3B, 2.7B, 6.7B, 13B, 30B).
Meta Evaluation To evaluate the alignment between our metrics and LM capabilities, we employed meta-evaluation by calculating the Spearman correlation coefficient between human-annotated ground truth benchmarks and our proposed refined informatics metrics. For the meta-evaluation experiments, we selected six benchmark datasets spanning four major domains as ground truth, corresponding to four key dimensions of large language model capabilities: Factuality: TruthfulQA Lin et al. (2022), FACTOR Muhlgay et al. (2024), Math: MATH Hendrycks et al. (2021), Reasoning: CommonsenseQA Talmor et al. (2019), TheoremQA Chen et al. (2023b), Knowledge: MMLU Hendrycks et al. (2020). We use the mean of all benchmark scores as the ground truth for the model’s comprehensive evaluation (CE).
@@
Baseline Metrics We selected purely representation-based baseline metrics that operate independently of ground-truth labels and model sampling, encompassing both informatics and geometric perspectives. The informatics metrics include Compression (DE) and Semantic Volume Li et al. (2025), while the geometric metrics consist of Curvature Hosseini and Fedorenko (2023) quantifying manifold curvature characteristics, and anisotropy. Diff-eRank Wei et al. (2024) is the metric that simultaneously models both information compression and geometric structure in language model representations, yet neglecting their direct synergistic relationship.
Baseline Anisotropy Razors In addition to PCS as the anisotropy razor for decoupling anisotropy from compression, we selected three anisotropy razors as baselines. Remove Directions Mu and Viswanath (2018) is a post-processing method for eliminating noisy directions. Whitening Su et al. (2021) eliminates correlations between features through global scaling, normalizing the eigenvalues to have the same mean and variance. LW Shrinkage Ledoit and Wolf (2004), on the other hand, adjusts extreme eigenvalues linearly towards the mean via Bayesian shrinkage.

### v3

# Compression Hacking: A Supplementary Perspective on Informatics Metric of Language Models from Geometric Distortion

@@ Abstract
Recently, the concept of “compression as intelligence” has provided a novel informatics metric perspective for language models (LMs), emphasizing that highly structured representations signify the intelligence level of LMs. However, from a geometric standpoint, the representation space of highly compressed LMs tends to degenerate into a highly anisotropic state, which hinders the LM’s ability to comprehend instructions and directly impacts its performance. We found this compression-anisotropy synchronicity is essentially the “Compression Hacking” in LM representations, where noise-dominated directions tend to create the illusion of high compression rates by sacrificing spatial uniformity. Based on this, we propose three refined compression metrics by incorporating geometric distortion analysis and integrate them into a self-evaluation pipeline. The refined metrics exhibit strong alignment with the LM’s comprehensive capabilities, achieving Spearman correlation coefficients above 0.9, significantly outperforming both the original compression and other internal structure-based metrics. This confirms that compression hacking substantially enhances the informatics interpretation of LMs by incorporating geometric distortion of representations  [^1] Codes: https://github.com/hggzjx/compression_hacking.
Compression Hacking: A Supplementary Perspective on Informatics Metric of Language Models from Geometric Distortion
Jianxiang Zang1, Meiling Ning2, Yongda Wei3, Shihan Dou1, Jiazheng Zhang1, Nijia Mo3, Binhong Li4, Tao Gui1thanks:  Corresponding author, Qi Zhang1, Xuanjing Huang1 1 Computation and Artificial Intelligence Innovative College, Fudan University, 2 Beijing University of Posts and Telecommunications 3 Shanghai University of International Business and Economics 4 Hong Kong University of Science and Technology (Guangzhou) jxzang25@m.fudan.edu.cn, tgui@fudan.edu.cn
@@ 1 Introduction
Recently, significant efforts have been devoted to exploring the mechanisms by which language models (LMs) process information internally, driving the development of LM self-evaluation wei2024diff; wang2024embedding; wang2024latent independent of specific tasks and model outputs. The concept of “compression as intelligence” sutskever2023compressors; deletanglanguage; chen2025information has provided a novel informatics interpretation for LMs, emphasizing that LMs eliminate redundant information through training while their representation spaces typically evolve from disordered to structured states. This property leads to a compression-based evaluation metric for LMs that utilizes differential entropy of representations, aiming to reflect model capabilities with their internal structural organization pichler2022differential; zhouyin2023understanding; li2025semantic. Existing studies have demonstrated strong alignment between this metric and LM scale wei2024diff; li2025semantic, which we have also empirically validated. However, as evidenced by the intuitive case where 175B GPT-3 brown2020language exhibits inferior overall capabilities compared to 32B Qwen2.5-Instruct hui2024qwen2, compression from a purely informatics standpoint, cannot fully align with LM capabilities, especially when comparing models from different families. Therefore, our research motivation is: Beyond information compression, what other properties should a metric quantify to effectively interpret the LMs’ intelligence level, and how should we model the relationships between these properties?
Relevant studies have shown that differences in model architecture and training paradigms inevitably lead to variations in the geometric structure of representations mimno2017strange; gao2019representation; skean2025layer. From a geometric standpoint, we were surprised to observe that LMs with high information compression tend to exhibit representation spaces that degenerate into highly anisotropic, distorted states. Highly anisotropic representations indicate varying sensitivity to semantic changes across different dimensions, which can hinder language models’ ability to comprehend instructions and consequently degrade their performance demeter2020stolen; yu2022rare; rudman2024stable.
@@
Using this framework, we evaluate 18 open-source LMs and conduct meta-evaluations on factuality, reasoning, math, and knowledge tasks to obtain ground-truth capability scores. Extensive experiments demonstrate that the refined metrics exhibit strong alignment with the LM’s comprehensive capabilities, achieving Spearman correlation coefficients above 0.9, which significantly outperforms both the original compression and other internal structure-based metrics. This validates that compression hacking substantially enhances the informatics interpretation of LMs by incorporating geometric distortion analysis of representations. The main contributions are summarized as follows:
@@
  • The refined metrics exhibit significantly stronger alignment with LM’s comprehensive capabilities compared to the original compression metric, thereby establishing a task-agnostic self-evaluation perspective for LMs.
@@ 2 Compression Hacking
In this section, we analyze the compression-anisotropy synchronicity in LM representations, where highly compressed LMs tend to exhibit in-context representations with strong anisotropy. Our investigation proceeds in two stages: First, we quantify both compression and anisotropy metrics by examining the internal structure of LM representations (covariance matrices). We then fit regression curves to model the relationship between anisotropy and compression, verifying its statistical significance. Second, through mechanistic analysis, we identify the underlying cause of this phenomenon, what we term “compression hacking”.
@@ 2.1 Preliminary: Differential Entropy based Compression Metric
The compression perspective provides an information-theoretic foundation for LM evaluation, revealing the intrinsic connections between model scale, generalization capability, and data volume, thus offering theoretical guidance for optimizing model design pichler2022differential; sutskever2023compressors; deletanglanguage; wei2024diff; chen2025information. Related studies have shown that the differential entropy $\mathcal{H}_{\text{DE}}(\mathbf{Z})=-\mathbb{E}_{\bm{w}\sim\mathcal{V}}\mathbf{z}(\bm{w})\log\mathbf{z}(\bm{w})$ of LM representations $\mathbf{z}(\bm{w})$ can reflect their compression capacity cheninside; zhouyin2023understanding; li2025semantic. Lower differential entropy suggests that the representations formed by nonlinear transformation, which removes redundant information, are closer to optimal coding. These representations exhibit more concentrated distributions and lower uncertainty, reflecting more efficient information compression deletang2023language. Semantic Volume leverages this property to model representation uncertainty li2025semantic.
We thus define compression metric as the negative differential entropy of representations (i.e., $\mathcal{C}_{\text{DE}}(\mathbf{Z})\overset{\rm def}{=}-\mathcal{H}_{\text{DE}}(\mathbf{Z})$). Since the differential entropy is equivalent to the logdet estimator cheninside of their covariance matrix, the compression metric follows the definition in Eq. 2.
$\mathcal{C}_{\text{DE}}(\mathbf{Z})$ $\overset{\rm def}{=}-\frac{1}{2}\text{logdet}\left(\Sigma_{\mathbf{Z}}\right)=-\frac{1}{2}\sum_{d=1}^{D}\log\lambda_{d}$ (2)
Figure 1: Comparison of compression metrics across different models and their corresponding ground-truth comprehensive capabilities, categorized into intra-family and cross-family comparisons.
We first conducted preliminary exploration to assess whether differential entropy-based compression metrics effectively reflect LM capabilities. Our evaluation included both intra-family (OPT family) and cross-family tests (Qwen2.5-3b-Instruct, Qwen2.5-7b, LLaMA3.1-8b, and OPT-13b), with ground-truth settings following Section 4.1. As shown in Figure 1, we found that compression metrics showed only positive correlations with model scale, consistent with related studies wei2024diff; li2025semantic. However, Figure 1(left) indicates that differential entropy-based compression is effective only for intra-family evaluation, while Figure 1(right) reveals its limited applicability across diverse architectures and training paradigms. These findings prompted our integration of geometric properties into compression analysis.
@@ 2.2 Anisotropy: The Geometric Property Correlated with Compression
The anisotropy of language models is a geometric property of representations that reflects the non-uniform distribution of semantics across different directions in the representation space ethayarajh2019contextual; cai2021isotropy; demeter2020stolen. Highly anisotropic representations hinder LMs’ ability to comprehend instructions yu2022rare; rudman2024stable, directly impairing their overall capabilities. We performed principal component analysis to visualize the in-context representation spaces of the aforementioned four models. As shown in Figure 2, we made the intriguing observation that models with higher compression levels consistently exhibited greater unevenness in their dimensional distributions, namely, higher anisotropy. This suggests a potential synergistic relationship between compression and anisotropy. If we can quantify this relationship and confirm its statistical significance, it could provide valuable guidance for refining compression metrics.
Current tools for qualitatively and quantitatively analyzing the anisotropy of language models mainly rely on similarity computations of representations ethayarajh2019contextual; cai2021isotropy; rudman2022isoscore. However, what we need is an anisotropy metric that can establish a connection with entropy-based information compression. Relevant studies arora2016latent; mu2018all have shown that the anisotropy measure $\mathcal{A}$ is mathematically defined as formulated in Eq. 3. We aim to extend this measure to relate to the internal structure of representations (eigenvalues of the covariance matrix).
@@
where $\mathcal{Z}(\mathbf{c})=\sum_{\bm{w}\in\mathcal{V}}\exp\left(\mathbf{c}^{\top}\mathbf{z}(\bm{w})\right)$ is the original partition function should approximately be a constant for any unit vector $\mathbf{c}$. $\mathcal{A}$ is a number greater than 1, where larger values indicate stronger anisotropy in the representation space. Ideally, this value should be as close to 1 as possible. Considering that $\arg\max_{\|\mathbf{c}\|=1}\mathcal{Z}(\mathbf{c})$ and $\arg\min_{\|\mathbf{c}\|=1}\mathcal{Z}(\mathbf{c})$ do not have closed-form solutions, we attempt to approximate $\mathcal{Z}(\mathbf{c})$ via Taylor expansion as formulated in Eq. 4.
@@ 2.3 Systematic Analysis
Significance Analysis Next, we analyze the significance of compression hacking, which manifests as compression-anisotropy synchronicity. Based on the aforementioned metrics, we calculated the estimates of both compression and anisotropy for instruction representations across four LMs in our preliminary experiments, both of which can be exclusively represented by the eigenvalues of the representation covariance matrix. Given their characteristic patterns, we modeled a linear regression of compression against the logarithmic values of anisotropy, as shown in Figure 3. The regression analysis reveals two key findings through R² and p-values: (1) compression as the dependent variable can be well and significantly explained by anisotropy, and (2) Mann-Whitney U tests mcknight2010mann confirm statistically significant differences in regression curves across different models.
Figure 3: Regression fitting curves of compression versus anisotropy for different models, along with Mann-Whitney U tests between them. Here, **** denotes statistical significance at the 0.01% level.
@@ 3.1 Refined Metrics
Spectral Entropy Quantification Figure 2 illustrates that, from the perspective of eigenvalue distribution, the mechanism of compression hacking is that the secondary components causing anisotropy ($\lambda_{d}$) are homologous to the principal components of the compression part ($\log\lambda_{d}$). Interesting, spectral entropy roy2007effective precisely models this characteristic, and it is formally equivalent to a compression metric weighted by eigenvalues (Compression (SE)), as formulated in Eq. 8.
@@
Semantic Coefficient of Variation Just as compression-anisotropy synchronicity serves as a distinct manifestation of compression hacking, where compression is characterized by the mean of eigenvalue logarithms (reflecting the overall volume of the embedding space li2025semantic), while anisotropy corresponds to the ratio of extreme eigenvalues (quantifying the variation of semantic embeddings across different dimensions). Thus, we formulate their ratio as the Semantic Coefficient of Variation (Semantic CV) in Eq. 9. This metric accurately characterizes the magnitude of anisotropy relative to information compression in the representation space $\mathbf{Z}$.
@@
Manifold Correction Protocol Numerous studies have proposed train-free “anisotropy razors” to reduce the anisotropy of representation space in a train-free manner, thereby enhancing representational capacity mu2018all; su2021whitening. This inspires us to decouple anisotropy from compression by selecting an appropriate anisotropy razor. Considering the exponential sharp decline in the eigenvalues of principal components corresponding to preceding dimensions due to compression hacking, we propose Principal Component Smoothing (PCS) as an anisotropy razor, inspired by the LW-shrinkage ledoit2004well. By setting a smoothing coefficient $\beta\in[0,1]$ (default value is set to 0.9), we shift the representation space toward principal directions, resulting in a flatter transformed feature spectrum. This transformation is based on the covariance matrix of the representation and is achieved by defining the mapping $\mathcal{T}_{\text{PCS}}$ as formulated in Eq. 10, thereby refining the compression metric Compression (PCS). In Theorem B.2, we prove that under sparse spectrum conditions, the PCS estimator exhibits higher statistical stability than the LW shrinkage.
@@ 4.1 Setup
Models Since our evaluation focuses on the internal structure of model representations, we evaluated 18 open-source language models from three different model families with varying sizes. These families are the LLaMA3 family grattafiori2024llama (LLaMA3.2-1B, LLaMA3.2-1B-Instruct, LLaMA3.2-3B, LLaMA3.2-3B-Instruct, LLaMA3.1-8B, LLaMA3.1-8B-Instruct), Qwen2.5-Instruct family hui2024qwen2 (0.5B, 1.5B, 3B, 7B, 14B, 32B), OPT family zhang2022opt (0.125B, 1.3B, 2.7B, 6.7B, 13B, 30B).
Meta Evaluation To evaluate the alignment between our metrics and LM capabilities, we employed meta-evaluation by calculating the Spearman correlation coefficient between human-annotated ground truth benchmarks and our proposed refined informatics metrics. For the meta-evaluation experiments, we selected six benchmark datasets spanning four major domains as ground truth, corresponding to four key dimensions of large language model capabilities: Factuality: TruthfulQA lin2022truthfulqa, FACTOR muhlgay2024generating, Math: MATH hendrycks2021measuring, Reasoning: CommonsenseQA talmor2019commonsenseqa, TheoremQA chen2023theoremqa, Knowledge: MMLU hendrycks2020measuring. We use the mean of all benchmark scores as the ground truth for the model’s comprehensive evaluation (CE).
@@
Baseline Metrics We selected purely representation-based baseline metrics that operate independently of ground-truth labels and model sampling, encompassing both informatics and geometric perspectives. The informatics metrics include Compression (DE) and Semantic Volume li2025semantic, while the geometric metrics consist of Curvature hosseini2023large quantifying manifold curvature characteristics, and anisotropy. Diff-eRank wei2024diff is the metric that simultaneously models both information compression and geometric structure in language model representations, yet neglecting their direct synergistic relationship.
Baseline Anisotropy Razors In addition to PCS as the anisotropy razor for decoupling anisotropy from compression, we selected three anisotropy razors as baselines. Remove Directions mu2018all is a post-processing method for eliminating noisy directions. Whitening su2021whitening eliminates correlations between features through global scaling, normalizing the eigenvalues to have the same mean and variance. LW Shrinkage ledoit2004well, on the other hand, adjusts extreme eigenvalues linearly towards the mean via Bayesian shrinkage.